# Channel Estimation for Beyond Diagonal RIS via Tensor Decomposition

### v1

@@ V-C Discussion
It is worth mentioning that the BTFK method accomplishes channel estimation in closed form utilizing $Q$ independent and parallel processing routines, which implies a shorter processing delay than BTALS. The latter consists of a sequential process of alternating estimation steps, where the channel matrices associated with the $Q$ BD-RIS groups are jointly estimated. Therefore, there is a longer processing delay than BTKF. From a practical perspective, optimized procedures for computing rank-one approximations and the possibility of parallel processing are attractive features of BTKF despite its higher training overhead than BTALS. Thus, there is clearly a trade-off between both algorithms involving complexity, training overhead, and processing delay.
Remark 3: The steps 4 and 5 of Algorithm 2, derived from the unfolding expressions (40)-(41), can be replaced by the updating steps $\bm{\hat{G}}_{(i)}=[\bm{\mathcal{Y}}]_{(1)}\big{[}[\bm{\mathcal{S}}]_{(1)}(\bm{I}_{K}\otimes\bm{H}_{(i-1)})^{\text{T}}\big{]}^{+}$ and $\bm{\hat{H}}_{(i)}=[\bm{\mathcal{Y}}]_{(2)}\big{[}[\bm{\mathcal{S}}]_{(2)}\left(\bm{I}_{K}\otimes\bm{G}_{(i)}\right)^{\text{T}}\big{]}^{+}$ that result from the pair of expressions (21)-(22). As mentioned in Remark 1, both representations produce the same result. Numerical experiments have shown us that the second choice yields a faster runtime in the MATLAB environment. This is because the traditional Kronecker product can be computed more efficiently than the block Kronecker product.
V-D Uniqueness and identifiability issues
We discuss the identifiability issues and requirements in terms of parameter settings related to an essentially unique estimate of the channel matrices using the proposed BTKF and BTALS algorithms. In particular, capitalizing on the uniqueness conditions for the BTD [24] allows us to derive useful relations involving the system parameters, such as the number of transmit/receive antennas and training length for the proposed algorithms.
@@ V-D Uniqueness and identifiability issues
First, let us recall that the received pilot tensor in the noise-free case follows a Tucker-$2$ decomposition, as shown in (20). In the group-connected case, it can be expressed as a sum of $Q$ blocks (see (28)). In the general case, Tucker models are not essentially unique due to rotational freedom since nonsingular transformations that compensate each other can be applied to the core tensor and the corresponding factor matrix in each mode without changing the tensor fit [30, 25]. Specifically, defining $\bm{T}_{\bm{G}}\in\mathbb{C}^{N\times N}$ and $\bm{T}_{\bm{H}}\in\mathbb{C}^{N\times N}$, any alternative solution $\hat{\bm{G}}=\bm{G}\bm{T}_{\bm{G}}$ and $\hat{\bm{H}}=\bm{H}\bm{T}_{\bm{H}}$ and $\hat{\bm{\mathcal{S}}}=\bm{\mathcal{S}}\times_{1}\bm{T}^{-1}_{\bm{G}}\times_{2}\bm{T}^{-1}_{\bm{H}}$ yields the same result since $\bm{\mathcal{Y}}=\hat{\bm{\mathcal{S}}}\times_{1}\hat{\bm{G}}\times_{2}\hat{\bm{H}}=(\bm{\mathcal{S}}\times_{1}\bm{T}^{-1}_{\bm{G}}\times_{2}\bm{T}^{-1}_{\bm{H}})\times_{1}(\bm{G}\bm{T}_{\bm{G}})\times_{2}(\bm{H}\bm{T}_{\bm{H}})=\bm{\mathcal{S}}\times_{1}(\bm{G}(\bm{T}_{\bm{G}}\bm{T}^{-1}_{\bm{G}}))\times_{2}(\bm{H}(\bm{T}_{\bm{H}}\bm{T}^{-1}_{\bm{H}}))=\bm{\mathcal{S}}\times_{1}\bm{G}\times_{2}\bm{H}$. In the group-connected case, these nonsingular transformation matrices are confined within each block, and one can also arbitrarily permute the $Q$ blocks without changing the result [24]. However, the core tensor is fixed in our case since we know the BD-RIS training tensor $\bm{\mathcal{S}}$ at the receiver. In this case, these transformation matrices reduce to trivial scalar indeterminacies affecting the factor matrices that compensate each other, i.e., $\bm{T}_{\bm{G}}=\alpha\bm{I}_{N}$ and $\bm{T}_{\bm{H}}=\beta\bm{I}_{N}$., with $\alpha\beta=1$.
Then, in the noise-free case, the estimated channel matrices $\hat{\bm{G}}$ and $\hat{\bm{H}}$ are related to the true ones by the identities
@@
Considering the 3-mode unfolding of the received pilot tensor given in (23) and introducing the transformation matrices, we have $\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}=[\bm{\mathcal{S}}]_{(3)}(\hat{\bm{H}}\otimes\hat{\bm{G}})^{\text{T}}=[\bm{\mathcal{S}}]_{(3)}((\bm{H}\bm{T}_{\bm{H}})\otimes(\bm{G}\bm{T}_{\bm{G}}))^{\text{T}}=[\bm{\mathcal{S}}]_{(3)}(\bm{T}_{\bm{H}}\otimes\bm{T}_{\bm{G}})^{\text{T}}(\bm{H}\otimes\bm{G})^{\text{T}}$. Note that the only choice for $\bm{T}_{\bm{G}}$ and $\bm{T}_{\bm{H}}$ satisfying the Kronecker-product equation $\bm{T}_{\bm{H}}\otimes\bm{T}_{\bm{G}}=\bm{I}_{N^{2}}$ are scaled identity matrices that compensate each other.
The reasoning for the group-connected case is similar but applies to each group. Hence, the channel matrices associated with the $q$-th group are related to the true ones by the identities
@@
for $q=1,\ldots,Q$. It is worth noting that these scaling ambiguities affecting the individual channels are irrelevant to the optimization of the BD-RIS scattering coefficients and transmit/receive beamforming vectors since they cancel each other when building the combined channel used to optimize the system parameters using state-of-the-art methods [5].
As discussed in Sections V-A and V-B (and summarized in Algorithms 1 and 2), the decoupled estimation of the channel matrices relies on the different unfoldings of the received pilot tensor. For the BTKF algorithm, the essential uniqueness of the channel estimates requires $\bar{\bm{S}}_{3}$ to be full row-rank so that the filtering step in (52) yields a unique solution. This implies that $K\geq\bar{N}^{2}Q$ is required. The BTALS algorithm estimates the channel matrices by solving the LS problems in (62) and (63) (see also steps 4 and 5 of Algorithm 2). Defining $\bm{P}_{1}\doteq\bm{S}_{1}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{H}\right)^{\text{T}}\in\mathbb{C}^{N\times KM_{T}}$ and $\bm{P}_{2}\doteq\bm{S}_{2}\left(\bm{\overline{I}}\hskip 1.42271pt|\otimes|\hskip 1.42271pt\bm{G}\right)^{\text{T}}\in\mathbb{C}^{N\times KM_{R}}$ and from the LS problems in (62) and (63), we conclude that identifiability of $\hat{\bm{G}}=[\bm{\mathcal{Y}}]_{(1)}\bm{P}_{1}^{\dagger}$ and $\hat{\bm{H}}=[\bm{\mathcal{Y}}]_{(2)}\bm{P}_{2}^{\dagger}$ requires that $\bm{P}_{1}$ and $\bm{P}_{2}$ be right-invertible, which implies $KM_{T}\geq N$ and $KM_{R}\geq N$, respectively. In addition, it is required that $\bm{S}_{1}\in\mathbb{C}^{\bar{N}Q\times\bar{N}KQ}$ and $\bm{S}_{2}\in\mathbb{C}^{\bar{N}Q\times\bar{N}KQ}$ have full row-rank. Note, however, that these conditions are necessary but not sufficient for uniqueness. Nevertheless, these conditions are useful for the system designer to eliminate “bad” configurations that do not yield unique estimates of $\bm{G}$ and $\bm{H}$ separately. Otherwise stated, system setups that violate one of these conditions do not lead to unique channel estimates and can be discarded.
Sufficient condition: We can obtain a condition for guaranteed identifiability of the channel matrices by adopting a similar reasoning as that of [24] to study the uniqueness of the block Tucker2 decomposition (therein referred to as type-2 decomposition). The discussion is simpler in our context since the core tensor (represented by the BD-RIS training tensor) is known. Under the assumption that $\bm{G}$ and $\bm{H}$ are full column-rank, their essential uniqueness is guaranteed if $K\geq 3$ and the BD-RIS training tensor $\bm{\mathcal{S}}$ does not have proportional frontal slices, which means that $\bm{\mathcal{S}}_{..k}\neq\alpha\bm{\mathcal{S}}_{..k^{\prime}}$, $\forall k\neq k^{\prime}$, $k,k^{\prime}\in\{1,\ldots,K\}$, where $\alpha$ is a scalar. This result implies that the design of the BD-RIS tensor is crucial to the uniqueness of the channel estimates for both BTKF and BTALS algorithms.
VI BD-RIS training tensor design
As shown in the previous section, the BTKF and BTALS algorithms exploit the BD-RIS training tensor differently to estimate the involved channel matrices. While the BTKF relies on its-mode unfolding (49) for a two-step channel separation using Kronecker factorization, BTALS exploits its 1-mode and 2-mode unfoldings for a direct and iterative channel estimation method. It is clear that a proper design of the BD-RIS training tensor is crucial to ensuring unique estimates of the channel matrices (up to trivial scaling ambiguities) for both algorithms. We are interested in a more flexible (lower overhead) design, which is valid for $K<<\bar{N}^{2}Q$, in contrast to the design of [5] that requires $K=\bar{N}^{2}Q$ for the LS estimation. It also allows us to go from conventional RIS to BD-RIS fully-connected architectures while covering a broader range of values of $K$. In the following, we discuss the design of the BD-RIS training tensor in two steps. We first formulate the baseline BD-RIS tensor design and then discuss the BD-RIS training structures used by the BTKF and BTALS algorithms, respectively.
@@ VI BD-RIS training tensor design
Let us define the tensor $\bm{\mathcal{Z}}\in\mathbb{C}^{\bar{N}\times\bar{N}\times K_{1}}$ such that its $k_{1}$-th frontal slice corresponds to
$\bm{\mathcal{Z}}_{..k_{1}}\doteq\textrm{unvec}_{\bar{N}\times\bar{N}}(\bm{z}_{k_{1}})\in\mathbb{C}^{\bar{N}\times\bar{N}},\,\,k_{1}=1,\ldots,K_{1}.$ (64)
Note that these frontal slices are column/row-permuted versions of the DFT matrix obtained from the applied circular shifts, as detailed earlier in this section. The basic structure of the BD-RIS tensor $\bm{\mathcal{S}}^{(q)}\in\mathbb{C}^{\bar{N}\times\bar{N}\times K}$ is given, as a function of its $k$-th frontal slice, as follows
@@
where $\bm{\Theta}\doteq[\bm{\theta}_{1},\ldots,\bm{\theta}_{Q}]\in\mathbb{C}^{K_{2}\times Q}$ is a (possibly truncated) $K_{2}\times Q$ Hadamard ($\bm{\Omega}_{\textrm{HAD}}$) or a DFT ($\bm{\Omega}_{\textrm{DFT}}$) matrix, where $K_{2}=K/K_{1}$ must be an integer. Note that each slice $\bm{\mathcal{S}}^{(q)}_{..k}\in\mathbb{C}^{N\times N}$ is orthogonal, i.e., $\bm{\mathcal{S}}^{(q)}_{..k}\bm{\mathcal{S}}^{(q)\text{H}}_{..k}=\bm{\mathcal{S}}^{(q)\text{H}}_{..k}\bm{\mathcal{S}}^{(q)}_{..k}=\bm{I}_{\bar{N}}$, $k=1,\ldots,K$ since $\bm{\mathcal{Z}}_{..k_{1}}$ defined in (64) is orthogonal $\forall k_{1}\in\{1,\ldots,K_{1}\}$. Collecting the $K$ slices in (65) according to (29)-(31) gives $[\bm{\mathcal{S}}^{(q)}]_{(1)}=\bm{\theta}^{\text{T}}_{q}\otimes[\bm{\mathcal{Z}}]_{(1)}\in\mathbb{C}^{\bar{N}\times\bar{N}K}$, $[\bm{\mathcal{S}}^{(q)}]_{(2)}=\bm{\theta}^{\text{T}}_{q}\otimes[\bm{\mathcal{Z}}]_{(2)}\in\mathbb{C}^{\bar{N}\times\bar{N}K}$, and $[\bm{\mathcal{S}}^{(q)}]_{(3)}=\bm{\theta}_{q}\otimes[\bm{\mathcal{Z}}]_{(3)}\in\mathbb{C}^{K\times\bar{N}^{2}}$.
Substituting into (42)-(43), and (49), we obtain

### v2

@@ V-B Block Tucker alternating least squares (BTALS) algorithm for decoupled channel estimation
Remark 3: Steps 4 and 5 of Algorithm 2 are based on the compact expressions for the 1-mode and 2-mode unfoldings in (57). In the fully connected case, we can alternatively consider the general expressions (21)-(22) and replace these updating steps with the $\bm{\hat{G}}_{(i)}=[\bm{\mathcal{Y}}]_{(1)}\big{[}[\bm{\mathcal{S}}]_{(1)}(\bm{I}_{K}\otimes\bm{H}_{(i-1)})^{\text{T}}\big{]}^{+}$ and $\bm{\hat{H}}_{(i)}=[\bm{\mathcal{Y}}]_{(2)}\big{[}[\bm{\mathcal{S}}]_{(2)}\left(\bm{I}_{K}\otimes\bm{G}_{(i)}\right)^{\text{T}}\big{]}^{+}$, respectively, to estimate the channel matrices directly.
V-C Uniqueness and scaling ambiguities
This section discusses the uniqueness issues related to the estimation of the channel matrices $\bm{G}$ and $\bm{H}$ from the Tucker tensor model using the BTKF and BTALS algorithms. In the following, we obtain applicable conditions involving the training parameters that ensure unique estimates of the channel matrices. First, note that the two algorithms exploit the structure of the received signal tensor differently. While BTKF is based on its 3-mode unfolding, BTALS uses its 1-mode and 2-mode unfoldings. This translates into different requirements involving the system parameters used for channel training.
Result 1: Assuming $\textrm{rank}(\bm{S}_{3})=\bar{N}^{2}Q$, unique solutions for the estimates of the channel matrices $\bm{G}$ and $\bm{H}$ are obtained using the BTKF algorithm according to problem (48).
Proof: Note that left-filtering $[\bm{\mathcal{Y}}]_{(3)}$ in (47) using the 3-mode unfolding of the BD-RIS training tensor $\bm{S}_{3}$ yields a unique estimate of the composite channel $\bm{H}\hskip 2.0pt\hskip 1.42271pt|\otimes|\hskip 1.42271pt\hskip 2.0pt\bm{G}$ if $\bm{S}_{3}$ has full column-rank (which requires $K\geq\bar{N}^{2}Q$). Under this condition, an additional constraint is necessary to extract the individual estimates of the channel matrices, which is accomplished from each block of the estimated composite channel from rank-one matrix approximation steps. $\blacksquare$
Result 2: Let us assume $\textrm{rank}(\bm{S}_{1})=\textrm{rank}(\bm{S}_{2})=\bar{N}Q$. Uniqueness of the estimates of the channel matrices $\bm{G}$ and $\bm{H}$ according to problems (60)-(61) using the BTALS algorithm requires $KM_{T}\geq\bar{N}Q$ and $KM_{R}\geq\bar{N}Q$.
Proof: Let us consider the LS estimation steps in (62). Defining $\bm{P}_{1}\doteq\bm{T}_{\bm{H}}\bm{S}_{1}^{\text{T}}\in\mathbb{C}^{KM_{T}\times\bar{N}Q}$ and $\bm{P}_{2}\doteq\bm{T}_{\bm{G}}\bm{S}_{2}^{\text{T}}\in\mathbb{C}^{KM_{R}\times\bar{N}Q}$, the uniqueness of $\hat{\bm{G}}=[\bm{\mathcal{Y}}]_{(1)}[\bm{P}_{1}^{\text{T}}]^{\dagger}$ and $\hat{\bm{H}}=[\bm{\mathcal{Y}}]_{(2)}[\bm{P}_{2}^{\text{T}}]^{\dagger}$ requires that $\textrm{rank}(\bm{P}_{1})=\textrm{rank}(\bm{P}_{2})=N$, so that $\bm{P}_{1}$ and $\bm{P}_{2}$ left-invertible, which implies $KM_{T}\geq N$ and $KM_{R}\geq N$. Note, however, that these conditions are necessary but not sufficient for uniqueness. Nevertheless, these conditions are useful for the system designer to eliminate invalid system setups that do not yield unique estimates of the channel matrices. Otherwise stated, system setups that violate one of these conditions can be discarded. $\blacksquare$
@@ V-C Uniqueness and scaling ambiguities
Under the conditions discussed previously, the estimates of $\bm{G}$ and $\bm{H}$ are unique up to scaling ambiguities. To verify this, recall that the received pilot tensor in the noise-free case follows a Tucker-$2$ decomposition (see (20)). Generically, Tucker models are not essentially unique due to rotational freedom since nonsingular transformations compensating each other can be applied to the factor matrix and the core tensor [38, 32]. Defining non-singular matrices $\bm{T}_{\bm{G}}\in\mathbb{C}^{N\times N}$ and $\bm{T}_{\bm{H}}\in\mathbb{C}^{N\times N}$, any alternative solution $\hat{\bm{G}}=\bm{G}\bm{T}_{\bm{G}}$, $\hat{\bm{H}}=\bm{H}\bm{T}_{\bm{H}}$ and $\hat{\bm{\mathcal{S}}}=\bm{\mathcal{S}}\times_{1}\bm{T}^{-1}_{\bm{G}}\times_{2}\bm{T}^{-1}_{\bm{H}}$ yields the same result since $\bm{\mathcal{Y}}=\hat{\bm{\mathcal{S}}}\times_{1}\hat{\bm{G}}\times_{2}\hat{\bm{H}}=(\bm{\mathcal{S}}\times_{1}\bm{T}^{-1}_{\bm{G}}\times_{2}\bm{T}^{-1}_{\bm{H}})\times_{1}(\bm{G}\bm{T}_{\bm{G}})\times_{2}(\bm{H}\bm{T}_{\bm{H}})=\bm{\mathcal{S}}\times_{1}(\bm{G}(\bm{T}_{\bm{G}}\bm{T}^{-1}_{\bm{G}}))\times_{2}(\bm{H}(\bm{T}_{\bm{H}}\bm{T}^{-1}_{\bm{H}}))=\bm{\mathcal{S}}\times_{1}\bm{G}\times_{2}\bm{H}$. In the group-connected case, these nonsingular transformation matrices are confined within each block, and one can also arbitrarily permute the $Q$ blocks without changing the result [31]. However, the core tensor is fixed since the BD-RIS training tensor $\bm{\mathcal{S}}$ is known at the receiver. In this case, the only feasible structures for $\bm{T}_{\bm{G}}$ and $\bm{T}_{\bm{H}}$ correspond to scaled identity matrices that compensate each other, i.e., $\bm{T}_{\bm{G}}=\alpha\bm{I}_{N}$ and $\bm{T}_{\bm{H}}=\beta\bm{I}_{N}$., with $\alpha\beta=1$. Thus, any alternative solutions for $\hat{\bm{G}}$ and $\hat{\bm{H}}$ satisfy the identities
@@
These identities can alternatively be understood by looking at the 3-mode unfolding in (23) and introducing the transformation matrices, yielding $\big{[}\bm{\mathcal{Y}}\big{]}_{\left(3\right)}=[\bm{\mathcal{S}}]_{(3)}(\hat{\bm{H}}\otimes\hat{\bm{G}})^{\text{T}}=[\bm{\mathcal{S}}]_{(3)}((\bm{H}\bm{T}_{\bm{H}})\otimes(\bm{G}\bm{T}_{\bm{G}}))^{\text{T}}=[\bm{\mathcal{S}}]_{(3)}(\bm{T}_{\bm{H}}\otimes\bm{T}_{\bm{G}})^{\text{T}}(\bm{H}\otimes\bm{G})^{\text{T}}$. Note that the only choice for $\bm{T}_{\bm{G}}$ and $\bm{T}_{\bm{H}}$ satisfying the Kronecker-product equation $\bm{T}_{\bm{H}}\otimes\bm{T}_{\bm{G}}=\bm{I}_{N^{2}}$ are scaled identity matrices that compensate each other.
The reasoning for the group-connected case is similar, but instead applies to each group. Defining the block-diagonal transformation matrices $\bm{T}_{G}=\textrm{blkdiag}(\bm{T}^{(1)}_{G},\ldots,\bm{T}^{(Q)}_{G})$ and $\bm{T}_{H}=\textrm{blkdiag}(\bm{T}^{(1)}_{H},\ldots,\bm{T}^{(Q)}_{H})$ that act independently over each channel block, and from similar arguments, alternative channel matrices $\hat{\bm{G}}^{(q)}$ and $\hat{\bm{H}}^{(q)}$, associated with the $q$-th BD-RIS group, satisfy the following identities
@@
for $q=1,\ldots,Q$. Hence, individual channel estimation involves a single scalar ambiguity in the fully connected case and $Q$ scalar ambiguities in the group-connected case.
@@
It is worth noting, however, that these scaling ambiguities affecting the individual channels are irrelevant to the optimization of the BD-RIS scattering coefficients and transmit/receive beamforming vectors since they cancel each other when building the composite channel used to optimize the system parameters using state-of-the-art methods [5].
VI BD-RIS training design
As previously discussed, the BTKF and BTALS algorithms exploit the BD-RIS training tensor differently to estimate the individual channel matrices. While BTKF relies on its 3-mode unfolding (45) for a two-step channel separation using Kronecker factorization, BTALS exploits its 1-mode and 2-mode unfoldings in (57) for an iterative estimation of the channel matrices using alternating least squares. Proper design of the BD-RIS tensor is crucial to obtaining reasonable estimates of the individual channel matrices using these algorithms.
To address the physical constraints of group-connected BD-RIS, the work [5] proposed an orthogonal design that requires $K=\bar{N}^{2}Q$. We propose a BD-RIS training design with a more flexible construction, which covers a broader range of values of $K$, while being valid for smaller training lengths $K<\bar{N}^{2}Q$, as opposed to the design of [5].
@@ VI BD-RIS training design
Let $\bm{Z}\doteq[\bm{z}_{1},\ldots,\bm{z}_{K_{1}}]\in\mathbb{C}^{\bar{N}^{2}\times K_{1}}$ be as a (possibly truncated) circulant matrix constructed such that $\bm{z}_{k_{1}}\doteq\textrm{circshift}(\textrm{vec}(\bm{\Omega}_{\textrm{DFT}}),k_{1}-1)$, $k_{1}=1,\ldots,K_{1}$, where $K_{1}=\textrm{min}(K,\bar{N}^{2})$, and $\bm{\Omega}_{\textrm{DFT}}\in\mathbb{C}^{\bar{N}\times\bar{N}}$ is a DFT matrix, with $\bm{\Omega}^{\textrm{H}}_{\textrm{DFT}}\bm{\Omega}_{\textrm{DFT}}=\bm{I}_{\bar{N}}$. Define the tensor $\bm{\mathcal{Z}}\in\mathbb{C}^{\bar{N}\times\bar{N}\times K_{1}}$ whose $k_{1}$-th frontal slice corresponds to
$\bm{\mathcal{Z}}_{..k_{1}}\doteq\textrm{unvec}_{\bar{N}\times\bar{N}}(\bm{z}_{k_{1}})\in\mathbb{C}^{\bar{N}\times\bar{N}},\,\,k_{1}=1,\ldots,K_{1},$ (64)
which are column/row-permuted versions of $\bm{\Omega}_{\textrm{DFT}}$. The basic structure of $\bm{\mathcal{S}}^{(q)}\in\mathbb{C}^{\bar{N}\times\bar{N}\times K}$ has frontal slices given by
@@
where $\bm{\Theta}\doteq[\bm{\theta}_{1},\ldots,\bm{\theta}_{Q}]\in\mathbb{C}^{K_{2}\times Q}$ is a (possibly truncated) $K_{2}\times Q$ Hadamard or a DFT matrix, where $K_{2}=K/K_{1}$ must be an integer. Collecting the $K$ slices in (65) according to (29)-(31) gives $[\bm{\mathcal{S}}^{(q)}]_{(1)}=\bm{\theta}^{\text{T}}_{q}\otimes[\bm{\mathcal{Z}}]_{(1)}\in\mathbb{C}^{\bar{N}\times\bar{N}K}$, $[\bm{\mathcal{S}}^{(q)}]_{(2)}=\bm{\theta}^{\text{T}}_{q}\otimes[\bm{\mathcal{Z}}]_{(2)}\in\mathbb{C}^{\bar{N}\times\bar{N}K}$, and $[\bm{\mathcal{S}}^{(q)}]_{(3)}=\bm{\theta}_{q}\otimes[\bm{\mathcal{Z}}]_{(3)}\in\mathbb{C}^{K\times\bar{N}^{2}}$, from which we obtain